# XAI-CF – Examining the Role of Explainable Artificial Intelligence in Cyber Forensics

### v2

@@ 5.4.1 Trustworthy
An XAI-CF system is used to make important and critical decisions, especially in a court of law. Before making such decisions based on the results and pieces of evidence provided by an XAI-CF system stakeholders of CF must trust the XAI-CF system. Therefore, this is one of the most critical prerequisites of an XAI-CF system. Trust in an XAI-CF system influences a user’s ability to believe in the output of the system without any uncertainty (bhatt2020explainable, ). The foremost property that can increase the trust of a user in an XAI-CF system is the quality of explanations provided by the system. Generally, users trust explanations created by humans rather than machines (kunkel2019let, ). An explanation by a machine can be improved if we integrate human-in-the-loop in an XAI-CF system. In this way, both of them complement each other and can produce more trustworthy results. This integration is still in its infancy and is one of the challenges we discuss later in the paper. A human-meaningful explanation by an XAI-CF system also increases trust in the system (nourani2019effects, ). Here the human-meaningful explanation means understandable by humans. As there are different stakeholders of CF the explanation should be tailored to each group of stakeholders. The trust in an XAI-CF system increases as the user becomes more familiar and experienced with the system over its time of use (holliday2016user, ; yang2017evaluating, ). For an XAI-CF system to become trustworthy this familiarity and experience should be positive. The trust of the user on an XAI-CF system can be measured using interview questions, both closed and open-ended, and also lickert-scale (albaum1997likert, ) questionnaires can be used to measure the user’s trust. Transparency (felzmann2019transparency, ) is another property of an XAI-CF system that increases trust. Making information open and transparent requires that the user should be literate enough to make use of the information provided and can assess the risks of XAI-CF decision-making systems. Different stakeholders of CF have different abilities to make use of the information provided, therefore, transparency should be tailored to the user’s ability. For example, an XAI-CF system should be transparent enough to a forensics analyst so that he can extract the right pieces of evidence to help in making an informed decision. Similarly, the forensics analyst should be familiar enough of the XAI-CF system to complete such a task.
@@ 5.4.2 Understandable
An XAI-CF system should be understandable to the stakeholders of CF. Interpretability is the ability to explain a model in understandable terms to humans (doshi2017towards, ). Understandability and interpretability are complementary. Clarity and Comprehensibility are the two other properties that an explanation should have to be more understandable, and all these three are also complementary to understandability. Understandability depends on the perception of the user. The explanations provided by an XAI-CF system help a user build a mental model of how the system works. In cognitive psychology mental models (johnson1983mental, ) gives a person the ability of reasoning and decision making. Visualization is one of the best approaches to explaining a system and easiest to understand. One of the methods that can be used to assist humans in building a better mental model of an XAI-CF system is to represent the system, including explanations, graphically/visually, such as causal loop, stock, and flow diagrams (groesser2012mental, ). Causal loop diagrams (CLDs) (haraldsson2004introduction, ) can be used to explicitly map an XAI-CF system and make it transparent and visible to others. A CLD consists of three basic elements, boxes, connections, and feedback loops. CLDs aid in analyzing a system qualitatively. To perform a quantitative analysis a CLD is transformed to a stock and flow diagram (sterman2001system, ). All these visualization techniques are being used in a wide range of areas to understand the system dynamics and we recommend using such techniques to help increase the understandability of an XAI-CF system. There is not much work done in building better mental models to increase the understandability of an XAI system. This area is open to a lot of future work. For example, such an XAI-CF system can use boxes/nodes to represent input features and output results, connections/edges can represent the correlation or influence of input features to output results, and feedback loops can represent different iterations in the system. We can use tags/labels both at nodes and edges to improve the flow of information for better understanding. In the future, we intend to work in this area and explore this idea further by designing and developing such an XAI-CF system.
@@ 5.4.3 Authentic
[…]
@@ 5.4.4 Interactive
[…]
@@ 6 Literature Review
This section discusses and reviews the current works on the application of XAI in CF. We divide the work into two classes: Statistical – works that use statistical techniques to interpret or explain results of AI CF methods, and Machine Learning (ML) – Works that use ML techniques to interpret or explain results of AI CF methods. In the following, we first present a summary of the literature review by using three tables, and then discuss and review the articles in detail.
Table 5 lists the articles reviewed in this section with the classification and the publication years. Seven articles are classified under statistical and were published between the years 2014 – 2023, whereas 12 articles are classified under ML and were published between the years 2018 – 2023.
Table 5: Classification of the XAI-CF research works reviewed in section 6
Classification
Articles
Publication Years
Statistical
(timmer2014extracting, ), (verheij2014catch, ), (vlek2014building, ), (timmer2015structure, ), (vlek2016method, ), (verheij2016arguments, ), (gopinath2023explainable, )
2014 – 2023
Machine Learning
(khan2018eliciting, ), (aditya2018enabling, ), (afzaliseresht2019explainable, ), (costantini2020assessing, ), (bolle2020role, ), (mahajan2021explainable, ), (nguyen2021human, ), (pethe2022atle2fc, ), (hall2022proof, ), (jayakumar2022visually, ), (de2022interpretable, ), (bouter2023protoexplorer, )
2018 – 2023
@@
Table 6 lists the different approaches employed (out of those mentioned in section 5.2) by the research works reviewed in this section. Most of the works employed the approach feature influence which shows its popularity and ease of implementation. The approaches Model simplification and visualization are employed in the same proportion, and seems like they are also popular but difficult to implement. Human-in-the-loop approach is least employed because it is a very new phenomenon but hopefully will be more popular and employed in the future.
Table 6: Approach employed (out of those mentioned in section 5.2) by the XAI-CF research works reviewed in section 6
Approaches
Articles
Feature Influence
(timmer2014extracting, ), (vlek2016method, ), (gopinath2023explainable, ), (khan2018eliciting, ), (aditya2018enabling, ), (costantini2020assessing, ), (bolle2020role, ), (mahajan2021explainable, ), (pethe2022atle2fc, ), (hall2022proof, )
Model Simplification
(verheij2014catch, ), (vlek2014building, ), (verheij2016arguments, ), (aditya2018enabling, ), (afzaliseresht2019explainable, ), (de2022interpretable, )
Visualization
(timmer2014extracting, ), (timmer2015structure, ), (jayakumar2022visually, ), (bouter2023protoexplorer, )
Human-In-The-Loop
(nguyen2021human, ), (bouter2023protoexplorer, )
@@
A summary of the articles reviewed in this section is shown in Table 7. The research works reviewed are published in the last ten years (2014 – 2023). Although the word XAI was coined in 2018 (gunning2019darpa, ) there are research works carried out before 2018 on explaining the statistical (part of AI) pieces of evidence presented in a legal case which is part of the CF as per definition 2 in this paper. Only two works use the state-of-the-art XAI method LIME to evaluate and explain AI models. Bayesian networks are popular for visualizing legal pieces of evidence and reasoning. Most of the statistical techniques are used to explain these Bayesian networks for the legal stakeholders. Out of these nineteen research works reviewed here the most modern and advanced approach employed is presented in (bouter2023protoexplorer, ) in the year 2023, by visualizing and using a human-in-the-loop to interact and improve the XAI model with each iteration. This shows that the current recent work in this area is moving in the right direction. This effort is appreciated but relatively rare compared to other XAI applications in other domains. Therefore, we must promote and facilitate this process to keep pace with the ever-changing technology in the present digital age.
@@
Article Year Types
Summary
@@
(timmer2014extracting, ) 2014 G S
Extract legal arguments from Bayesian networks using ranked inference rules and visualize them using argument graphs
(verheij2014catch, ) 2014 G S
Combines arguments, scenarios, and Bayesian network to formalize the reasoning with legal pieces of evidence
(vlek2014building, ) 2014 G S
Combines arguments, extended scenarios, and Bayesian network to formalize the reasoning with legal pieces of evidence
(timmer2015structure, ) 2015 S S
Extract legal arguments from Bayesian networks using ranked inference rules and use Markov blanket to build support graphs to reduce the redundant arguments
(vlek2016method, ) 2016 G S
@@
Extend (vlek2014building, ) by providing a probabilistic interpretation of scenario quality for explaining and understanding Bayesian networks with scenarios
(verheij2016arguments, ) 2016 G S
@@
Present a hybrid model that connects arguments and scenarios to improve coherent scenario-based interpretations that can assist in explanation
(khan2018eliciting, ) 2018 G A
@@
Associative rule mining to analyze security event logs and an action-based model is proposed for automatic examination for investigation purposes
(aditya2018enabling, ) 2018 G S
@@
Adversary testing framework to develop trust in a black box deep neural networks models in the context of CF
(afzaliseresht2019explainable, ) 2018 G A
@@
A contextualized interpretation of security events that can help explain the events and hence reduce the efforts of humans to interpret events
(costantini2020assessing, ) 2020 I A
@@
A framework that provides transparency to the reasoning process of evidence collection during forensic analysis of IoT devices for investigation purposes.
(bolle2020role, ) 2020 G A
@@
A framework based on scientific interpretation principles and formalizing different steps such as forensics evaluation to help evaluate, understand, and explain the outputs of automated AI CF system
(mahajan2021explainable, ) 2021 G A
@@
Apply LIME, one of the XAI techniques, to evaluate AI models for forensics analysis of toxic comments
(nguyen2021human, ) 2021 I A
@@
Integrates human (forensics expert) and machine in an interactive and iterative cycle that uses human intelligence to drive the XAI-CF system
(pethe2022atle2fc, ) 2022 I S
@@
A model with an explainable layer that consists of Frequent Pattern Growth, one of the XAI techniques, and a deep learning classifier
(hall2022proof, ) 2022 G A
@@
Apply LIME, one of the XAI techniques, to evaluate AI models for explaining forensics analysis of image and video file content
(jayakumar2022visually, ) 2022 G S
@@
Apply Anchors, one of the XAI techniques, to provide visual explanations for deep fake detection for improving the interpretability of the model
(de2022interpretable, ) 2022 N A
@@
A federated transformer log learning model that provides forensic investigators with actionable information to effectively analyze system activity and operation
(gopinath2023explainable, ) 2023 I A
@@
Present metrics, such as the steps and time consumed in image creation and data retrieval, that can increase trust in the pieces of evidence in IoT devices
(bouter2023protoexplorer, ) 2023 G S
@@
The forensics analyst can explore and refine the visualized model and provide an explanation of how the model aided in making the decision
@@ 6.1 Statistical
Statistics plays a vital role in AI methods (friedrich2022there, ). This section presents a review and discussion of some of the research works that explain AI systems in CF using statistical techniques. Although there are some research works, such as (hepler2007object, ; keppens2012argument, ; fenton2013general, ) done before 2014, we do not include them in the literature review because of the constraints of the published date, and also some of the later works (2014 and latter) are extensions of these works. Generally, three main approaches are used to formalize reasoning about legal evidence. The narrative approach is also called scenarios, argumentative approach, and probabilistic (usually Bayesian networks) approach. All the works discussed here either use one or combine two or more of these approaches to formalize a legal case for reasoning about the pieces of evidence.
Timmer et al. (timmer2014extracting, ) apply inference rules to extract arguments from forensic Bayesian networks and these inferences are then displayed (visualized) as arguments graphs. For each rule, its strength (influence on the evidence) is computed that are based on the conditional probabilities. Rules are ranked according to their strength. These rules are then combined into arguments understandable by a legal expert. This improves the communication between legal and forensic experts. The paper cites examples of perversion of justice due to misunderstandings in probabilistic reasoning. Statistical methods used in the paper address the specific challenges associated with reasoning under uncertainty in legal cases involving forensic evidence. It helps bridge the gap between probabilistic and argumentative reasoning and supports decision-making while mitigating the risk of miscarriages of justice.
Verheij et al. (verheij2014catch, ) combine arguments, scenarios, and probabilities (Bayesian network) to formalize the reasoning with legal pieces of evidence. Both qualitative and quantitative information can be broken down into what is relevant in the legal case. Scenarios here are the events and tell us about what has happened. The evidence supports, by providing arguments, one or more scenarios, and this support can be specified both qualitatively and quantitatively. Additional evidence can change the support of a scenario. By collecting more pieces of evidence a scenario support can change from possibility to certainty. The authors use logic to formalize the process of integrating arguments, scenarios, and probabilities. This formalization improves the communication between legal and forensic experts.
Vlek et al. (vlek2014building, ) also combine arguments, scenarios, and probabilities (Bayesian network) to formalize the reasoning with legal pieces of evidence. The paper addresses the problem of how to represent and evaluate existing scenarios generated from an investigation using probabilistic techniques. To represent a crime scenario, the list of idioms used in (fenton2013general, ) is extended with four narrative idioms. The authors present heuristics to select relevant scenarios. A scenario can also be unfolded to provide more details to give more insight into the legal evidence. A Bayesian network is built out of these scenarios that can be used for reasoning about legal pieces of evidence. The development of idioms for legal arguments within the context of Bayesian networks aims to provide a structured and coherent approach. In the context of XAI-CF, making the reasoning and inference transparent and providing explanations for why certain conclusions are reached, is crucial in legal and decision-making contexts. Also, It can aid in establishing causal relationships within the Bayesian networks, allowing for a better understanding of how evidence and arguments are linked and supporting causal reasoning, which is often central to legal cases.
Timmer et al. (timmer2015structure, ) extended their previous work (timmer2014extracting, ) and reduced the number of redundant argumentative results. This is achieved by grouping the useful arguments in one cluster using a Markov blanket (pearl1988probabilistic, ). The Markov blanket is then used to build a support graph, The support graph model provides different ways of probabilistic reasoning about legal pieces of evidence. The Markov blanket, being a statistical concept, is used to improve the efficiency of argumentation models. In the context of XAI-CF, it can be important to explain how Markov blankets are used to optimize the probabilistic reasoning process, helping legal experts and decision-makers understand the statistical principles at play.
@@
Vlek et al. (vlek2016method, ) proposed a method for explaining and understanding Bayesian networks with scenarios. The authors adapted a previously proposed scheme (vlek2014building, ) for building a Bayesian network with scenarios. Scenarios in text form can be extracted from a Bayesian network and presented in a report with evidential support and information about the quality of the scenario. This makes the Bayesian network understandable to a legal expert. The paper mentions scenario quality and evidential support, indicating a need for understanding the strength of the presented scenarios. The quality of scenarios can be contributed by being explained in terms of probabilistic interpretations, and insights can be provided into how well each scenario is supported by the available evidence.
@@
Verheij et al. (verheij2016arguments, ) present a study of connections between arguments, scenarios, and probabilities and how they affect reasoning with pieces of evidence. In this study, an overview of combining the three approaches in different ways, such as pairwise and all three is presented and reviewed. This study helps in understanding the connections between probabilistic and non-probabilistic pieces of evidence that are useful in criminal investigation and decision-making. The authors claim that this will reduce reasoning errors and miscommunication between legal and forensics experts. The goal of the study is to reduce miscommunication between legal and forensic experts. Transparent and clear explanations of probabilistic information can play a role in achieving this goal. Also, the authors mention a hybrid model that connects arguments and scenarios. The incorporation of probabilistic elements into this hybrid model and the bridging between adversarial argumentative approaches and globally coherent scenario-based interpretations can be assisted in explanation.
@@
Gopinath et al. (gopinath2023explainable, ) perform forensic analysis of IoT devices in a method that increases trust in the analysis. They perform investigations on digital evidence and present certain metrics to increase trust in the pieces of evidence. These metrics are reports of different scenarios in detail, such as steps to retrieve the data, how the report was generated, etc, and the time consumed in image creation and data retrieval. These metrics are obtained using statistical methods for a comprehensive understanding and application of quantitative methods in these scenarios. These metrics require little effort on the part of the forensic analyst but can greatly help in the decision process of whether to trust the pieces of evidence collected or not.

### v3

@@ 5.4.1 Trustworthy
An XAI-CF system is used to make important and critical decisions, especially in a court of law. Before making such decisions based on the results and pieces of evidence provided by an XAI-CF system, stakeholders of CF must trust the XAI-CF system. Therefore, this is one of the most critical prerequisites of an XAI-CF system. Trust in an XAI-CF system influences a user’s ability to believe in the output of the system without any uncertainty [178]. The foremost property that can increase the trust of a user in an XAI-CF system is the quality of explanations provided by the system. Generally, users trust explanations created by humans rather than machines [179]. An explanation by a machine can be improved if we integrate human-in-the-loop in an XAI-CF system. In this way, both of them complement each other and can produce more trustworthy results. This integration is still in its infancy and is one of the challenges we discuss later in the paper. A human-meaningful explanation by an XAI-CF system also increases trust in the system [180]. Here, the human-meaningful explanation means understandable by humans. As there are different stakeholders of CF, the explanation should be tailored to each group of stakeholders. The trust in an XAI-CF system increases as the user becomes more familiar and experienced with the system over its time of use [181, 182]. For an XAI-CF system to become trustworthy, this familiarity and experience should be positive. The trust of the user on an XAI-CF system can be measured using interview questions, both closed and open-ended, and also Likert-scale [183] questionnaires can be used to measure the user’s trust. Transparency [184] is another property of an XAI-CF system that increases trust. Making information open and transparent requires that the user be literate enough to make use of the information provided and can assess the risks of XAI-CF decision-making systems. Different stakeholders of CF have different abilities to make use of the information provided; therefore, transparency should be tailored to the user’s ability. For example, an XAI-CF system should be transparent enough to a forensics analyst so that they can extract the right pieces of evidence to help in making an informed decision. Similarly, the forensics analyst should be familiar enough with the XAI-CF system to complete such a task.
@@ 5.4.2 Understandable
An XAI-CF system should be understandable to the stakeholders of CF. Interpretability is the ability to explain a model in understandable terms to humans [185]. Understandability and interpretability are complementary. Clarity and Comprehensibility are the two other properties that an explanation should have to be more understandable, and all three are also complementary to understandability. Understandability depends on the perception of the user. The explanations provided by an XAI-CF system help a user build a mental model of how the system works. In cognitive psychology, mental models [186] give a person the ability to reason and make decisions. Visualization is one of the best approaches to explaining a system and easiest to understand. One of the methods that can be used to assist humans in building a better mental model of an XAI-CF system is to represent the system, including explanations, graphically/visually, such as causal loop, stock, and flow diagrams [187]. Causal loop diagrams (CLDs) [188] can be used to explicitly map an XAI-CF system and make it transparent and visible to others. A CLD consists of three basic elements: boxes, connections, and feedback loops. CLDs aid in analyzing a system qualitatively. To perform a quantitative analysis, a CLD is transformed to a stock and flow diagram [189]. All these visualization techniques are being used in a wide range of areas to understand the system dynamics and we recommend using such techniques to help increase the understandability of an XAI-CF system. There is not much work done in building better mental models to increase the understandability of an XAI system. This area is open to a lot of future work. For example, such an XAI-CF system can use boxes/nodes to represent input features and output results, connections/edges can represent the correlation or influence of input features to output results, and feedback loops can represent different iterations in the system. We can use tags/labels both at nodes and edges to improve the flow of information for better understanding. In the future, we intend to work in this area and explore this idea further by designing and developing such an XAI-CF system.
@@ 5.4.3 Authentic
An explanation is authentic when it accurately renders and expresses a model’s behavior [190]. To understand a model we need to know how the model behaves under different conditions. The explanation of this behavior needs to be rendered correctly without any bias. Another property that makes an XAI-CF system authentic is completeness. There is the famous quote “Justice is blind”, but as we know, sources of information can bias the judicial process [191, 192]. The pieces of evidence produced by an XAI-CF system should be unbiased and complete. Bias can occur in an AI system for different reasons: a human-in-the-loop in an AI system; the inability to generate training data that is truly representative of the entire population of an AI system, etc. It is one of the major challenges of AI to keep the system unbiased. It may not be possible to remove all the biases; therefore, whenever a piece of evidence is presented that is generated by an XAI-CF, we should explain the biases in the system for the stakeholders to make an informed decision. Authenticity also requires an XAI-CF system to adhere to the legal compliance guidelines to make the information admissible in legal proceedings. For example, we should pay attention to the legal rules surrounding the collection and use of digital evidence; otherwise, it makes the evidence worthless and can leave the investigators vulnerable to countersuits. A piece of forensic evidence must satisfy two main conditions: it must be relevant [193]; and derived by scientific methods supported by appropriate validation [194]. For more details and coverage of the legal aspects of CF, interested readers are referred to [195]. One of the most critical rules of CF is not to change the original data. While collecting the evidence, forensic analysts use write blockers to follow this rule. Write blocking preserves the original evidence by preventing any data from being written to the original evidence device [106]. This ensures the collection of a genuine evidence. For an XAI-CF system to be authentic, it needs to preserve the genuineness of the pieces of evidence.
@@ 5.4.4 Interactive
After an initial explanation is presented, a user may have follow-up questions/queries [30]. To let the user perform such an action, we need to allow the user to interact with the XAI-CF system. For example, in a forensic deep fake analysis of a video [157], a forensics analyst zooms in on a part of the video image to observe the contributions of the XAI-CF system to the prediction score, to decide if the video is manipulated or not. Arya et al. [196] distinguish between static and interactive explanations. The first one does not change in response to feedback from the users, whereas the latter allows users to interact and drill down or ask questions until they are satisfied. XUI (Explanation User Interface) is a new concept and is defined as “the sum of outputs of an XAI system that the user can directly interact with” [197]. There are two modes of XUI [198], explanatory, where it gives a single explanation (i.e., visual or text), exploratory, where it lets users freely explore the AI model behavior. In this paper, we recommend the interactive explanation provided by the exploratory XUI as defined above. For an XAI-CF system to be interactive, it needs an easy-to-use Human Computer Interface (HCI). One of the basic goals of HCI is to improve the visual design for the interaction of users. Usability [199] can be applied to improve the visual design of an XAI-CF system. This is one of the major factors that affect the satisfaction of a user. Another important requirement for a good interactive system is user-centered-design [200]. This means that the system is designed based on an explicit understanding of the users, tasks, and environment. This is an iterative design process where in each iteration, the focus is on the users and their needs. A grand challenge of HCI is the design of interfaces that “allow users to understand the underlying computational processes better” [201]. There is limited research reported on HCI and XAI [202, 198] and very few [157] on HCI and XAI-CF. To build XAI-CF systems and make them successful in practice, there is a strong need to further research in this area and provide the user with a usable interactive XAI-CF system.
@@ 6 Literature Review
This section discusses and reviews the current works on the application of XAI in CF. We divide the work into two classes: Statistical – works that use statistical techniques to interpret or explain results of AI CF methods, and Machine Learning (ML) – Works that use ML techniques to interpret or explain results of AI CF methods. Figure 10 presents a complete taxonomy of the works that are reviewed in this section.
Figure 10: The taxonomy of the works discussed in section 6. Some of the works overlap, because they cover more than one taxonomy.
@@
Three articles are classified under statistical and were published between the years 2016 – 2023, whereas twelve articles are classified under ML and were published between the years 2018 – 2025. Most of the works employed the approach feature influence, which shows its popularity and ease of implementation. The approach Model simplification is also employed almost in the same proportion as the feature influence, and it seems like it is also popular but difficult to implement. visualization and Human-in-the-loop approaches are the least employed because they are very new phenomena and difficult to implement, but hopefully will be more popular and employed in the future.
@@
A summary of the articles reviewed in this section is shown in Table 5. The research works reviewed are published in the last ten years (2016 – 2025). Although the word XAI was coined in 2018 [177], there are research works carried out before 2018 on explaining the statistical (part of AI) pieces of evidence presented in a legal case, which is part of the CF as per definition 2 in this paper. Only two works use the state-of-the-art XAI method LIME to evaluate and explain AI models. Bayesian networks are popular for visualizing legal pieces of evidence and reasoning. Most of the statistical techniques are used to explain these Bayesian networks for the legal stakeholders. Out of these nineteen research works reviewed here, the most modern and advanced approaches employed are presented in [157] and [203] in the years 2023 and 2025. [157] used visualization and a human-in-the-loop to interact and improve the XAI model with each iteration. [203] presents a holistic framework to integrate XAI across the CF pipeline. This shows that the current recent work in this area is moving in the right direction. This effort is appreciated but relatively rare compared to other XAI applications in other domains. Therefore, we must promote and facilitate this process to keep pace with the ever-changing technology in the present digital age.
@@
Article Year Types Summary
@@
[147] 2016 G S Extend [144] by providing a probabilistic interpretation of scenario quality for explaining and understanding Bayesian networks with scenarios
@@
[146] 2016 G S Present a hybrid model that connects arguments and scenarios to improve coherent scenario-based interpretations that can assist in explanation
@@
[204] 2018 G A Associative rule mining to analyze security event logs and an action-based model is proposed for automatic examination for investigation purposes
@@
[148] 2018 G S Adversary testing framework to develop trust in a black box deep neural networks models in the context of CF
@@
[149] 2019 G A A contextualized interpretation of security events that can help explain the events and hence reduce the efforts of humans to interpret events
@@
[150] 2020 I A A framework that provides transparency to the reasoning process of evidence collection during forensic analysis of IoT devices for investigation purposes.
@@
[151] 2020 G A A framework based on scientific interpretation principles and formalizing different steps such as forensics evaluation to help evaluate, understand, and explain the outputs of automated AI CF system
@@
[152] 2021 G A Apply LIME, one of the XAI techniques, to evaluate AI models for forensics analysis of toxic comments
@@
[159] 2021 I A Integrates human (forensics expert) and machine in an interactive and iterative cycle that uses human intelligence to drive the XAI-CF system
@@
[153] 2022 I S A model with an explainable layer that consists of Frequent Pattern Growth, one of the XAI techniques, and a deep learning classifier
@@
[154] 2022 G A Apply LIME, one of the XAI techniques, to evaluate AI models for explaining forensics analysis of image and video file content
@@
[155] 2022 G S Apply Anchors, one of the XAI techniques, to provide visual explanations for deep fake detection for improving the interpretability of the model
@@
[156] 2022 N A A federated transformer log learning model that provides forensic investigators with actionable information to effectively analyze system activity and operation
@@
[205] 2023 I A Present metrics, such as the steps and time consumed in image creation and data retrieval, that can increase trust in the pieces of evidence in IoT devices
@@
[157] 2023 G S The forensics analyst can explore and refine the visualized model and provide an explanation of how the model aided in making the decision
[206] 2024 G A Optimizing file-type identification with integration of XAI methods, LIME and SHAP, to show which regions or features led to the predicted file type.
[203] 2024 N A A holistic framework that embeds XAI techniques, such as SHAP, LIME, and Anchors, etc., throughout a CF workflow.
[207] 2025 N A Compares SHAP and LIME explanations when applied to two classifiers (XGBoost and TabNet) for network intrusion detection.
[208] 2025 G A Explores unsupervised learning paired with XAI techniques for forensic domains that produce large volumes of unlabeled data.
@@ 6.1 Statistical
Statistics plays a vital role in AI methods [209]. This section presents a review and discussion of some of the research works that explain AI systems in CF using statistical techniques. Although there are several research works, such as [139, 140, 141, 142, 143, 144, 145] done before 2016, we do not include them in the literature review because of the constraints of the published date.
@@
Vlek et al. [147] proposed a method for explaining and understanding Bayesian networks with scenarios. The authors adapted a previously proposed scheme [144] for building a Bayesian network with scenarios. Scenarios in text form can be extracted from a Bayesian network and presented in a report with evidential support and information about the quality of the scenario. This makes the Bayesian network understandable to a legal expert. The paper mentions scenario quality and evidential support, indicating a need for understanding the strength of the presented scenarios. The quality of scenarios can be contributed by being explained in terms of probabilistic interpretations, and insights can be provided into how well each scenario is supported by the available evidence.
@@
Verheij et al. [146] present a study of connections between arguments, scenarios, and probabilities and how they affect reasoning with pieces of evidence. In this study, an overview of combining the three approaches in different ways, such as pairwise and all three is presented and reviewed. This study helps in understanding the connections between probabilistic and non-probabilistic pieces of evidence that are useful in criminal investigation and decision-making. The authors claim that this will reduce reasoning errors and miscommunication between legal and forensics experts. The goal of the study is to reduce miscommunication between legal and forensic experts. Transparent and clear explanations of probabilistic information can play a role in achieving this goal. Also, the authors mention a hybrid model that connects arguments and scenarios. The incorporation of probabilistic elements into this hybrid model and the bridging between adversarial argumentative approaches and globally coherent scenario-based interpretations can be assisted in explanation.
@@
Gopinath et al. [205] perform forensic analysis of IoT devices in a method that increases trust in the analysis. They perform investigations on digital evidence and present certain metrics to increase trust in the pieces of evidence. These metrics are reports of different scenarios in detail, such as steps to retrieve the data, how the report was generated, etc, and the time consumed in image creation and data retrieval. These metrics are obtained using statistical methods for a comprehensive understanding and application of quantitative methods in these scenarios. These metrics require little effort on the part of the forensic analyst but can greatly help in the decision process of whether to trust the pieces of evidence collected or not.